# Rotary Masked Autoencoders are Versatile Learners

### v1

@@ C.2 UEA Multivariate Time-series Experimental Setup
We use the same pre-training setting across all datasets, shown in Table 13. When fine-tuning, we find it necessary to have per-dataset hyperparameters. These are shown in Table 12. Because the various datasets have varying sizes, resulting in vastly different numbers of training steps, we choose to represent the length of the learning rate warmup as a percentage of all steps.
C.3 Pendulum Dataset
@@
C.4 Spiral dataset
Hyperparameter Value
Model size Tiny
Optimizer AdamW
AdamW betas $\beta_{1}=0.9,\ \beta_{2}=0.999$
Weight Decay 0.01
Base LR $3\times 10^{-4}$
Batch size 32
Epochs 500
Gradient clip 1
Linear LR warmup steps 2000
LR schedule Cosine
Dropout 0
Stochastic depth $\lambda$ 0
Precision FP32
C.5 Absolute Position Reconstruction Hyperparameters
Table 17: End-to-end training hyperparameters for the absolute reconstruction range experiment (Section A.3). Values that are constant across all runs are reported only once.
@@ C.5 Absolute Position Reconstruction Hyperparameters
Here we provide full details for the experiments shown in Section 5.1 and Appendix A.3. These results can be seen in Table 17 and Table 18, respectively. Across both experiments we keep the model size equal to RoMAE-tiny as described in Table 7 except for $d_{\text{model}}$ which we set to 960 for the experiment in Section A.3, and set according to the corresponding model size for the experiment in Section 5.1. All experiments in Section 5.1 use the same hyperparameters shown in Table 18. For both experiments we report the mean and standard deviation across 5 different seeds.
C.6 ELAsTiCC2 Experimental Setup
Table 19: ELAsTiCC2 full training hyperparameters.
@@ C.6 ELAsTiCC2 Experimental Setup
Batch size 8192 1024
Epochs 400 25
@@
Precision BF16 BF16
Full pre-training and fine-tuning hyperparameters for Section 5.3 can be found in Table 19. When creating the train/test split we use the code and pre-processing provided in the ATAT [7] code release [^7] https://github.com/alercebroker/ATAT. Because the input values in ELAsTiCC2 are nearly always larger than what one would find with images (e.g., of the order of 10-50 as opposed to between 0 and 1 with images), we found it beneficial to increase the gradient clip threshold to 10 from the common value of 1. In order to handle the variable number of points per sample we utilize padding, applying a pad mask to the attention scores.
We experimented with both FP16 and BF16 for mixed precision training, and found that FP16 resulted in NaN values. This is likely due to to the increased input range in ELAsTiCC2 interacting poorly with the reduced range of FP16. We found that BF16, with its larger range, worked well.
The processed ELAsTiCC2 dataset contains both the flux difference and variance for each point. An example training sample is visualized in Figure 4. In the figure one can see how each band has a different number of points, each of which is at a different time than the others. To convert each point to an embedding we use a patch size of (1, 2) for time and flux/variance respectively. Therefore, during pre-training the model predicts not only the masked flux difference values but also variance for each point, while time is embedded using position.
Figure 4: A training example from the ELAsTiCC2 dataset. The flux difference of each band has already been normalized.
C.7 Synthetic dataset
@@
C.8 PhysioNet
We adopt the pre–processed release of the PhysioNet/CinC 2012 Challenge [50], comprising multivariate clinical time–series collected during the 48h window following intensive–care–unit (ICU) admission. Static covariates (Age, Gender, Height, ICU type) occupy feature indices 003333 and are always observed, whereas the remaining $37$ channels are sparsely and irregularly sampled.
@@ C.8 PhysioNet
The two–dimensional positional vector $p_{n}$ encodes (i) the normalised time $t/T\!\in\![0,1]$ and (ii) the feature index $d$, providing the $n_{\text{pos}}=2$ co-ordinates required by RoMAE. During training we stochastically subsample $50\%$ of the observed tokens. The final input tensor hence has length $N$ for the values, and shape $(2,N)$ for the positions, and a Boolean mask of length $N$ indicating which tokens RoMAE must reconstruct, exactly matching the interpolation protocol of the HeTVAE benchmark.

### v2

@@ D.3 UEA Multivariate Time-series Experimental Setup
We use the same pre-training setting across all datasets, shown in Table 13. When fine-tuning, we find it necessary to have per-dataset hyperparameters. These are shown in Table 12. Because the datasets have varying sizes, resulting in vastly different numbers of training steps, we choose to scale the number of learning rate warmup steps as a percentage of total steps.
D.4 Pendulum Dataset
@@
D.5 Absolute Position Reconstruction Hyperparameters
Table 16: End-to-end training hyperparameters for the absolute reconstruction range experiment (Section B.1). Values that are constant across all runs are reported only once.
@@ D.5 Absolute Position Reconstruction Hyperparameters
Here we provide full details for the experiments shown in Section 5.1 and Appendix B.1. These results can be seen in Table 16 and Table 17, respectively. Across both experiments we keep the model size equal to RoMAE-tiny as described in Table 8 except for $d_{\text{model}}$ which we set to 960 for the experiment in Section B.1, and set according to the corresponding model size for the experiment in Section 5.1. All experiments in Section 5.1 use the same hyperparameters shown in Table 17. For both experiments we report the mean and standard deviation across 5 different seeds.
D.6 ELAsTiCC Experimental Setup
Table 18: ELAsTiCC full training hyperparameters.
@@ D.6 ELAsTiCC Experimental Setup
Batch size 16384 4096
Epochs 200 25
@@
Precision FP32 FP32
Full pre-training and fine-tuning hyperparameters for Section 5.4 can be found in Table 18. When creating the train/test split we use the code and pre-processing provided in the ATAT [7] code release [^7] https://github.com/alercebroker/ATAT. Because the input values in ELAsTiCC are nearly always larger than what one would find with images (e.g., of the order of 10-50 as opposed to between 0 and 1 with images), we found it beneficial to increase the gradient clip threshold to 10 from the common value of 1. In order to handle the variable number of points per sample we utilize padding, applying a pad mask to the attention scores. Although our final model was trained using full FP32 precision, we tested RoMAE with both FP16 and BF16 for mixed precision training, and found that FP16 resulted in NaN values. This is likely due to to the increased input range in ELAsTiCC interacting poorly with the reduced range of FP16. We found that BF16, with its larger range, worked well.
Each light curve in the ELAsTiCC dataset contains recordings of both the flux difference and variance. An example training sample is visualized in Figure 6, showing how each band has a different number of observations, each of which is at a different time than the others. To convert each point in the light curve to an embedding we use a patch size of (1, 2) for time and flux/variance respectively. Therefore, during pre-training the model predicts not only the masked flux difference values but also variance for each point, while time and band index are embedded using position.
Figure 6: A training example from the ELAsTiCC dataset. The flux difference of each band has already been normalized.
D.7 Spiral dataset
Hyperparameter Value
Model size Tiny
Optimizer AdamW
AdamW betas $\beta_{1}=0.9,\ \beta_{2}=0.999$
Weight Decay 0.01
Base LR $3\times 10^{-4}$
Batch size 32
Epochs 500
Gradient clip 1
Linear LR warmup steps 2000
LR schedule Cosine
Dropout 0
Stochastic depth $\lambda$ 0
Precision FP32
D.8 Synthetic dataset
@@
D.9 PhysioNet
We adopt the pre–processed release of the PhysioNet/CinC 2012 Challenge [50], comprising multivariate clinical time–series collected during the 48h window following intensive–care–unit (ICU) admission. Static covariates (Age, Gender, Height, ICU type) occupy feature indices 033 and are always observed, whereas the remaining $37$ channels are sparsely and irregularly sampled.
@@ D.9 PhysioNet
The two–dimensional positional vector $p_{n}$ encodes (i) the normalised time $t/T\!\in\![0,1]$ and (ii) the feature index $d$, providing the $n_{\text{pos}}=2$ co‑ordinates required by RoMAE. During training we stochastically subsample $50\%$ of the observed tokens. The final input tensor hence has length $N$ for the values, and shape $(2,N)$ for the positions, and a Boolean mask of length $N$ indicating which tokens RoMAE must reconstruct, exactly matching the interpolation protocol of the HeTVAE benchmark.